# FrameFusion: Combining Similarity and Importance for Video Token Reduction on Large Visual Language Models

### v1

@@ 6 Conclusion
In this paper, we propose FrameFusion, a similarity-based token merging method for video LVLMs. By combining similarity-based merging with importance-based pruning, FrameFusion reduces redundant visual tokens while retaining critical information. This approach optimizes computational efficiency and memory usage, enabling accurate video understanding with significantly fewer tokens. Experiments across multiple benchmarks demonstrate a 70% reduction in vision tokens with minimal performance loss, achieving over 3.1$\times$ LLM speedups and 1.6$\times$ end-to-end speedups. FrameFusion providing new insights in token similarity for LVLMs, offering an efficient and scalable solution for real-world video language applications.
@@ 7 Detailed Experiment Setup
For FrameFusion, the merging ratios across layers are controlled by two hyperparameters: $S_{\text{threshold}}$ and $N_{\text{threshold}}$, as discussed in Section 4.1.
$S_{\text{threshold}}$ defines the minimum cosine similarity required for two tokens to be considered similar and merged. Since similarity distributions vary across models, we set $S_{\text{threshold}}$ to match the median of similarity at the first model layer under typical input cases, such as 128 samples from the VideoMME dataset. For the Llava-Video series, we set $S_{\text{threshold}}=0.6$; while for MiniCPM-V, we set $S_{\text{threshold}}=0.7$.
@@ 8 Additional Experiment Results
8.1 Performance Across Token Budgets
@@
8.2 Retrieval Analysis
8.3 Speedup Across Model Sizes
We evaluate the scalability of FrameFusion ’s speedup across different model sizes, as shown in Figure 13. To accommodate the increased KV-Cache and memory overhead, we distribute the models across multiple GPUs. With larger models, FrameFusion achieves greater end-to-end speedups, delivering 2.8×2.8$\times$2.8 × for Llava-Video-32B on two GPUs and 3.2×3.2$\times$3.2 × for Llava-Video-72B on four GPUs at a 30% token budget.
Figure 13: Runtime of (a) Llava-Video-32B on two A100-80GB GPUs and (b) Llava-Video-72B on four A100-80GB GPUs using FrameFusion. Numbers on bars indicate LLM and end-to-end speedups.
8.4 Token Reduction Details
8.5 Scalability Numeric Results
@@
Model Method Knowledge Film & Television Sports Competition Artistic Performance Life Record Multilingual
Llava-Video-7B Original 63.1 67.2 61.8 61.7 63.7 58.9
StreamingLLM 55.1 57.2 56.0 54.2 52.9 48.9
FastV 59.1 60.0 58.9 57.8 58.1 55.6
Ours 62.7 63.6 58.0 61.7 60.8 56.7
Llava-Video-32B Original 60.9 68.9 57.3 59.2 58.1 58.9
StreamingLLM 53.1 55.6 50.2 53.3 49.5 53.3
FastV 56.1 62.2 55.6 56.7 54.3 50.0
Ours 58.9 65.6 55.6 59.2 54.3 54.4
Llava-Video-72B Original 73.2 74.4 68.0 71.4 68.9 62.2
StreamingLLM 65.7 66.7 59.3 65.6 58.7 58.9
FastV 66.8 72.8 61.1 69.2 63.2 61.1
Ours 72.2 73.1 65.6 69.7 65.9 61.1
Table 6: Numeric VideoMME scores of different methods and model sizes across various video categories.
Table 6 provides the VideoMME scores for various model sizes across different video categories, offering a numerical breakdown of Figure 13. FrameFusion consistently outperforms the baseline methods across diverse model sizes and video categories.
@@ 9 Asymptotic Complexity Analysis
We estimate the computing cost of FrameFusion following the approach of FastV [2]. Given a model with $L$ layers and a specified relative token budget $C$, FrameFusion operates in the merging stage from layer 0 to layer $K-1$, then transitions to the pruning stage at layer $K$. Let $N_{l}$ denote the number of tokens layer $l$ before token reduction at this layer. Note that $N_{l+1}$ represents the number of tokens of layer $l$ after token reduction, and we let $N_{-1}$ equal the original input token length $N$. FrameFusion reduces $N_{l}$ with merging and pruning at the initial $K+1$ layers. After the token reduction, the remaining tokens for the successive layers are calculated as follows:
@@
Note that the additional computation $F^{\prime}$ introduced by FrameFusion shows neglectable asymptotic complexity with respect to input length and model size, comparing with the $O(N^{2}D)$ and $O(ND^{2})$ complexities of the original model.
@@ 10.1 Similarity Distribution Details
We take 128 videos from the VideoMME dataset and calculate the variance in token similarity across different layers. As shown in Figure 15, the similarity variance decreases in the deeper layers of the model, validating Observation 3.
@@ 10.2 Observations on Additional Models
In addition to the analysis of the Llava-Video model in Section 3, we conduct a similar study on the MiniCPM architecture. Results are presented in Figures 16, 17, 18, and 19.
Overall, the conclusions align with those of the Llava-Video model, with a few notable differences: Firstly, as shown in Figure 16, MiniCPM, which incorporates Q-Former [29, 15], exhibits additional high similarity among vision tokens within the same frame. However, the prominent 210th sub-diagonal persists, supporting our token similarity calculation strategy. Secondly, as shown in Figure 17, high similarity decreases less steeply in deeper layers for MiniCPM compared to Llava-Video. Despite this, the superior efficiency of cascaded merging at shallower layers ensures that Design Choice 4.2 remains valid.
@@
Figure 18: Spearman Rank Correlation between adjacent layers for the MiniCPM-V-8B model.
@@ 10.3 Video Pruning Visualization
We select a video example to visualize the effect of our token merging strategy. Figure 20 shows the frames of the original video sampled at a frame rate of 1 fps. In Figure 21, we present the video input to the model after token merging in Layer 0, where blank patches indicate tokens that have been merged. Furthermore, we replace the blank regions with the average of the merged patches, and the resulting visualization is shown in Figure 22. As shown in the examples, FrameFusion token merging strategy successfully merges similar vision tokens, reducing the computational costs, while maintaining high vadility of the video.
@@ 11 Limitation and Future Works
While FrameFusion demonstrates significant improvements in token reduction and efficiency for video LVLMs, certain challenges remain for future works. First, the similarity-based merging process can be further refined to better handle highly diverse or complex video content, minimizing potential information loss. Second, the reliance on pre-defined similarity and importance metrics calls for the development of adaptive and task-specific strategies to improve generalization across diverse scenarios. Future work will focus on designing more robust similarity measures and integrating FrameFusion with advanced token-efficient architectures.

### v2

@@ 6 Conclusion
In this paper, we propose FrameFusion, a similarity-based token merging method for video LVLMs. By combining similarity-based merging with importance-based pruning, FrameFusion reduces redundant visual tokens while retaining critical information. This approach optimizes computational efficiency and memory usage, enabling accurate video understanding with significantly fewer tokens. Experiments across multiple benchmarks demonstrate a 70% reduction in visual tokens with minimal performance loss, achieving 1.6-3.6$\times$ end-to-end speedups. FrameFusion providing new insights in token similarity for LVLMs, offering an efficient and scalable solution for real-world video language applications.
Acknowledgement
This work was supported by National Natural Science Foundation of China (No. 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164, 92364201), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist). We thank all the support from Infinigence-AI.
@@ 7.1.1 Baseline Setup
7.1.2 FrameFusion Setup
Figure 10: The workflow of FrameFusion when applied to LVLMs. At each layer, FrameFusion performs merging, pruning, or no action between the self-attention and feed-forward layers, depending on the current stage. The stage initially starts as “merge” and updates according to transition conditions.
Workflow details. For FrameFusion, token merging is only applied to visual tokens because they dominate input length and show higher similarity between adjacent frames, enabling $O(N)$ complexity merging. The detailed workflow of FrameFusion is shown in Figure 10.
Hyperparameters. The merging ratios across layers are controlled by two hyperparameters: $S_{\text{threshold}}$ and $N_{\text{threshold}}$, as discussed in Section 4.1.
$S_{\text{threshold}}$ defines the minimum cosine similarity required for two tokens to be considered similar and merged. Since similarity distributions vary across models, we set $S_{\text{threshold}}$ to match the median of similarity at the first model layer under typical input cases, such as 128 samples from the VideoMME dataset. For the Llava-Video series, we set $S_{\text{threshold}}=0.6$; for MiniCPM-V, we set $S_{\text{threshold}}=0.7$; for NVILA-2B,8B,15B, We set $S_{\text{threshold}}=0.6,0.75,0.8$, respectively.
@@ 7.1.2 FrameFusion Setup
7.2 Model Setup
We follow the default frame count settings for all models, except for NVILA-Lite-2B. Since NVILA-Lite-2B is not specifically trained for video tasks, we set its frame count to 64. For the Llava-Video series and Minicpm-V, the frame count is set to 64, while for NVILA-Video-8B and NVILA-Video-15B, it is set to 256.
@@ 8 Additional Experiment Results
8.1 Performance
8.1.1 Computation-Accuracy Trade-off
Figure 11: The accuracy-computation trade-offs of various token compression methods, tested on Llava-Video-7B with VideoMME benchmark. Original* represents the original model with reduced frame rates.
Figure 12: The accuracy-computation trade-offs of various token compression methods, tested on NVILA-8B with VideoNIAH benchmark. Original* represents the original model with reduced frame rates.
We further investigate the trade-off between computational cost and accuracy. We evaluate the Llava-Video-7B and NVILA-8B models on the VideoMME and VideoNIAH benchmark, respectively. The results are shown in Figure  11 and Figure  12. As the number of FLOPs decreases, other baseline methods exhibit a noticeable decline in accuracy, whereas FrameFusion maintains superior performance.
8.1.2 Performance Across Different Input Length
Figure 13: The VideoMME performances for the Llava-Video-7B across various numbers of input frames.
Figure 13 presents the performance of Llava-Video-7B on the VideoMME benchmark as the number of input frames varies from 8 to 128. Across all configurations, FrameFusion consistently outperforms the baseline methods, demonstrating its robustness to different input length.
8.1.3 Performance Across Different Token Budgets
@@
8.1.4 Performance Across Different Models
@@
Model Method Short Medium Long KL FT SC AP LR ML
Llava-Video-7B Original 75.8 61.7 52.2 63.1 67.2 61.8 61.7 63.7 58.9
StreamingLLM 63.4 54.1 46.4 55.1 57.2 56.0 54.2 52.9 48.9
FastV 68.4 58.0 49.6 59.1 60.0 58.9 57.8 58.1 55.6
PruMerge 69.7 60.1 50.2 59.1 63.6 59.1 58.9 60.6 57.8
Ours 74.0 59.8 50.0 62.7 63.6 58.0 61.7 60.8 56.7
Llava-Video-72B Original 80.9 69.7 62.1 73.2 74.4 68.0 71.4 68.9 62.2
StreamingLLM 68.2 59.9 59.8 65.7 66.7 59.3 65.6 58.7 58.9
FastV 73.0 64.9 60.2 66.8 72.8 61.1 69.2 63.2 61.1
PruMerge 74.0 65.8 60.3 70.4 73.6 62.9 68.3 61.6 54.4
Ours 78.3 67.9 60.9 72.2 73.1 65.6 69.7 65.9 61.1
NVILA-2B Original 61.4 48.9 42.4 47.2 56.4 49.8 55.0 51.0 52.2
StreamingLLM 52.9 43.9 40.3 43.0 49.2 46.2 50.6 44.4 43.3
FastV 53.7 45.6 40.8 43.8 49.7 46.2 51.4 46.7 43.3
PruMerge 53.9 45.0 43.1 43.5 52.8 45.8 51.7 48.1 45.6
Ours 61.3 47.0 43.0 48.3 55.6 48.2 55.3 49.8 45.6
NVILA-8B Original 74.9 62.1 54.7 64.8 66.4 62.2 61.9 63.7 63.3
StreamingLLM 61.2 53.8 48.0 54.9 57.8 54.7 52.5 52.2 55.6
FastV 72.0 56.7 50.0 60.7 62.8 57.6 57.8 58.9 57.8
PruMerge 67.6 54.9 48.3 57.3 61.1 56.0 54.7 56.2 55.6
Ours 74.2 57.7 51.3 60.7 65.3 59.3 58.6 62.1 58.9
NVILA-15B Original 77.3 64.7 55.3 67.2 68.1 62.7 63.3 66.2 66.7
StreamingLLM 63.8 57.4 54.3 60.6 60.6 55.6 58.6 56.5 60.0
FastV 69.2 58.7 53.9 62.8 63.3 57.1 60.8 58.1 63.3
PruMerge 66.0 59.3 52.6 61.0 61.1 55.1 57.5 59.8 61.1
Ours 73.2 62.3 55.0 64.6 68.1 60.9 61.1 62.5 65.6
MiniCPM-V-8B Original 69.1 56.6 49.8 59.0 63.6 54.2 63.3 54.9 60.0
StreamingLLM 61.1 51.8 48.4 54.6 58.1 52.2 56.4 49.7 55.6
FastV 67.1 53.9 49.2 57.2 59.2 53.8 60.8 54.6 56.7
Ours 69.7 54.1 48.3 57.9 63.1 53.8 60.3 54.4 56.7
Table 7: Numeric VideoMME scores of different methods and model sizes across various video categories. “KL”, “FT”, “SC”, “AP”, “LR”, “ML” are short for “Knowledge”, “Film & Television”, “Sports Competition”, “Artistic Performance”, “Life Record”, and “Multilingual”.
As shown in Figure 14, FrameFusion consistently outperforms FastV baseline across all model sizes and VideoMME categories, demonstrating comparative performance with the original model at a 30% relative token budget. Note that the model Llava-Video-32B has been removed by its author team. However, in order to demonstrate the generalization capability of our FrameFusion method across variable model sizes, we still include this model in the performance and efficiency tests here.
Table 7 provides the VideoMME scores for various model sizes across different video lengths and categories, offering a numerical breakdown of Figure 14.
@@
Setting Qwen2-VL-7B InternVL-2.5-8B
Original Ours Original Ours
w/o sub 55.9 58.4 63.1 62.3
w/sub 60.6 61.1 66.3 64.2
Table 9: Performance comparison between original and Framefusion of Qwen2-VL-7B and InternVL-2.5-8B on VideoMME.
We further test the performance of two extra models: Qwen2-VL-7B and InternVL-2.5-8B. As shown in Table 9, our method performs well compared to original models on VideoMME.
8.1.5 Retrieval Benchmark Details
8.1.6 Performance on Image Benchmark
We further investigate our method’s performance on an image benchmark: MMMU-Pro-standard. As shown in Table 10, although our method is not designed for image inputs, it still demonstrates comparative performance.
Model Original Fastv Ours
NVILA-2B 23.6 23.8 23.1
NVILA-8B 30.3 28.7 29.0
NVILA-15B 36.1 30.6 32.8
Table 10: The MMMU-Pro-standard performance across NVILA-2B, 8B, and 15B for different methods.
8.2 Efficiency
8.2.1 Efficiency Across Different Model Sizes
We evaluate the scalability of FrameFusion ’s efficiency across different model sizes, as shown in Figure 17 and 18. To accommodate the increased KV-Cache and memory overhead, we distribute models across multiple GPUs. With larger models, FrameFusion achieves greater end-to-end speedups, delivering 2.8×2.8$\times$2.8 × for Llava-Video-32B on two GPUs and 3.2×3.2$\times$3.2 × for Llava-Video-72B on four GPUs at a 30% token budget. Besides, FrameFusion reduces memory consumption for KV-Cache to 37% for Llava-Video-32B and 51% for Llava-Video-72B with a 30% token budget.
Figure 17: Runtime and memory breakdown of Llava-Video-32B on two A100-80GB GPUs using FrameFusion. A relative token budget of 1.0 represents the original dense model. Numbers on bars show (a) LLM and end-to-end speedups and (b) LLM’s KV-Cache and total relative memory.
Figure 18: Runtime and memory breakdown of Llava-Video-72B on four A100-80GB GPUs using FrameFusion. A relative token budget of 1.0 represents the original dense model. Numbers on bars show (a) LLM and end-to-end speedups and (b) LLM’s KV-Cache and total relative memory.
8.2.2 Token Reduction Details
8.3 Ablation Study
8.3.1 Similarity Computation Strategy
Figure 20: The average token similarity of the merged tokens for the first layer of Llava-Video-7B model across various merging rates.
We empirically study whether our approach successfully finds the most similar token pairs. All three $O(N)$ complexity strategies are compared against the posterior optimal upper bound, which merges the most similar tokens using the full $N\times N$ similarity computation. As shown in Figure 20, given different merging rate, the token pairs found by our method constantly shows the highest average similarity. We successfully reach 90% average similarity with only $1/10^{4}$ computing overhead. Further ablations are detailed in Section 5.5
8.3.2 Distance Metrics
FrameFusion adopts cosine similarity as the distance metric between tokens. To evaluate the impact of different distance metrics, we replace cosine similarity with the inner product, Minkowski-2, and Minkowski-1 distance. We test the performance of FrameFusion at a 30% token budget. As shown in Table 11, the average accuracy using cosine similarity is 2.8%, 1.4%, and 1.4% higher than the baseline metrics, respectively.
Choice VideoNIAH VideoMME NExt-QA Avg.
inner product 71.3 58.9 55.0 61.7
minkowski-2 71.3 60.9 57.1 63.1
minkowski-1 71.3 61.0 57.0 63.1
cosine similarity 75.1 61.4 56.9 64.5
Table 11: Performance of different distance calculation strategies with the same relative token budget of 30% on VideoNIAH, VideoMME, and NExt-QA.
8.3.3 Choice of Similarity Threshold
We conduct ablation studies on the sensitivity of the similarity ($S_{\text{threshold}}$) and merging-pruning transition ($N_{\text{threshold}}$) thresholds on NVILA-8B. As shown in Table 12, our method shows robust performance to threshold variations.
Threshold Value VideoNIAH VideoMME NeXT-QA-mc
$S_{\text{threshold}}$ 0.6 73.6 56.9 56.5
0.7 (default) 73.3 57.4 56.3
0.8 72.9 57.6 56.5
0.9 72.2 57.7 56.3
$N_{\text{threshold}}$ 0.1 (default) 73.3 57.4 56.3
0.2 74.0 57.6 56.5
0.3 74.2 57.5 56.5
Table 12: Performance of different similarity and merging-pruning transition thresholds on VideoNIAH, VideoMME, and NExt-QA.
8.3.4 Effect of Positional Embedding
We investigate the impact of positional embeddings on token similarity. Specifically, we compare models with and without positional embedding at the first layer and analyze the resulting changes in the similarity of the input hidden states to the second layer. The results show that the L1-norm of the similarity matrix changes by an absolute amount of $0.0087\pm 0.0010$, corresponding to a relative change of $2.73\%\pm 0.66\%$. It shows that the token contents, rather than the positional embeddings, dominate token similarity.
@@ 9 Asymptotic Complexity Analysis
We estimate the computing cost of FrameFusion following the approach of FastV [3]. Given a model with $L$ layers and a specified relative token budget $C$, FrameFusion operates in the merging stage from layer 0 to layer $K-1$, then transitions to the pruning stage at layer $K$. Let $N_{l}$ denote the number of tokens in layer $l$ before token reduction at this layer. Note that $N_{l+1}$ represents the number of tokens of layer $l$ after token reduction, and we let $N_{-1}$ equal the original input token length $N$. FrameFusion reduces $N_{l}$ with merging and pruning at the initial $K+1$ layers. After the token reduction, the remaining tokens for the successive layers are calculated as follows:
@@
Note that the additional computation $F^{\prime}$ introduced by FrameFusion shows negligible asymptotic complexity with respect to input length and model size, compared with the $O(N^{2}D)$ and $O(ND^{2})$ complexities of the original model.
@@ 10.1 Similarity Distribution Details
We take 128 videos from the VideoMME dataset and calculate the variance in token similarity across different layers. As shown in Figure 21, the similarity variance decreases in the deeper layers of the model, validating Observation 3. No significant outliers are observed in token similarity, in contrast to the common outliers seen with respect to the magnitude of hidden features [38, 36].
@@ 10.2 Observations on Additional Models
In addition to the analysis of the Llava-Video model in Section 3, we conduct a similar study on the MiniCPM architecture. Results are presented in Figures 22, 23, 24, and 25.
Overall, the conclusions align with those of the Llava-Video model, with a few notable differences: Firstly, as shown in Figure 22, MiniCPM, which incorporates Q-Former [34, 17], exhibits additional high similarity among visual tokens within the same frame. However, the prominent 210th sub-diagonal persists, supporting our token similarity calculation strategy. Secondly, as shown in Figure 23, high similarity decreases less steeply in deeper layers for MiniCPM compared to Llava-Video. Despite this, the superior efficiency of cascaded merging at shallower layers ensures that Design Choice 4.2 remains valid.
@@
Figure 24: Spearman Rank Correlation (SRC) between adjacent layers for the MiniCPM-V-8B model.
@@ 10.3 Video Pruning Visualization
We select a video example to visualize the effect of our token merging strategy. Figure 26 shows the frames of the original video sampled at a frame rate of 1 fps. In Figure 27, we present the video input to the model after token merging in Layer 0, where blank patches indicate tokens that have been merged. Furthermore, we replace the blank regions with the average of the merged patches, and the resulting visualization is shown in Figure 28. As shown in the examples, FrameFusion token merging strategy successfully merges similar visual tokens, reducing the computational costs, while maintaining high validity of the video.
@@ 10.4 Importance-Similarity Joint-Distribution
11 Additional Discussion on Related Works
Prior works have also explored token merging in image-based tasks [13, 45, 25]. For instance, while FrameFusion adopts an $O(N)$ temporal merging strategy, EVLGen [13] performs $O(N^{2})$ spatial merging via bipartite matching among tokens. AIM [45] similarly adopts bipartite-matching-based merging prior to the first layer of the LLM, followed by a token pruning process in subsequent LLM layers, ultimately reducing the number of visual tokens to zero. LLaVA-Prumerge [25] first prunes tokens at the output of the visual encoder and then merges the pruned tokens into the top-$k$ most similar remaining tokens. In all these methods, the similarity computation incurs a complexity of $O(N^{2})$. Although the computational efficiency of is comparable at the image scale ($N\approx 256$), our method scales more effectively to video scenarios where $N$ can reach 10K to 1M tokens.
@@ 12 Limitation and Future Works
While FrameFusion demonstrates significant improvements in token reduction and efficiency for video LVLMs, certain challenges remain for future work. First, the similarity-based merging process can be further refined to better handle highly diverse or complex video content, minimizing potential information loss. Second, the reliance on pre-defined similarity and importance metrics calls for the development of adaptive and task-specific strategies to improve generalization across diverse scenarios. Future work will focus on designing more robust similarity measures and integrating FrameFusion with advanced token-efficient architectures.